OpenStudy (anonymous):

Anyone know how to solve the summation Sum[1/Sqrt[2^i], {i, 0, -1 + k}] In other words the sum of 1/sqrt(2^i) from i=0 to k-1 ? Thanks!

OpenStudy (anonymous):

\[\begin{align*} \sum_{i=0}^{k-1}\frac{1}{\sqrt{2^i}}&=\frac{1}{\sqrt{2^0}}+\frac{1}{\sqrt{2^1}}+\frac{1}{\sqrt{2^2}}+\cdots+\frac{1}{\sqrt{2^{k-2}}}+\frac{1}{\sqrt{2^{k-1}}}\\ &=1+\frac{1}{\sqrt2}+\frac{1}{2}+\cdots+ \frac{1}{\sqrt{\frac{2^k}{4}}}+\frac{1}{\sqrt{\frac{2^k}{2}}}\\ &=1+\frac{1}{\sqrt2}+\frac{1}{2}+\cdots+ \frac{2}{\sqrt{2^k}}+\frac{\sqrt2}{\sqrt{2^k}}\\ &=\frac{1}{\sqrt{2^k}}\left(\sqrt{2^k}+\frac{\sqrt{2^k}}{\sqrt2}+\frac{\sqrt{2^k}}{2}+\cdots+2+\sqrt2\right) \end{align*} \] Hmm, I'm not sure what to do now, but according to WolframAlpha, I seem to be headed in the right direction with factoring out that 1/√(2^k): http://www.wolframalpha.com/input/?i=sum+1%2Fsqrt%282%5Ei%29+from+i%3D0+to+i%3Dk-1

OpenStudy (anonymous):

Thanks for the help!